OpenStudy (anonymous):

What is the period of the function y=2sinx

OpenStudy (anonymous):

Ah the period of the function \(f(x)=2\sin(x)\) is the same as the function \(g(x)=\sin(x)\). That would be the \(\sin(x)\) repeats every \(2\pi\)